OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)? 1 2 -2 -1

OpenStudy (gorv):

(-4-1)/(3-(-2))=-5/5=-1

OpenStudy (gorv):

@kwashondapearson97